OpenStudy (anonymous):

what is themass of s sample of water that absorbed 10,000j of heat energy, causing its temperature to rise from 35.1 degrees to 48.6 degrees?

OpenStudy (frostbite):

Hello @lilyliu9276 and welcome to \(\LARGE \sf \bbox[#40B9E9]{\color{white}{Open}}\bbox[#A8CE91]{\color{white}{Study}}\) In this question we are going to be using the equation that relates the energy of a sample to the temperature with the heat capacity as constant. We are going to be using that \[\Large Q=m \times c \times \Delta T\] Where \(Q\) is the heat added, \(m\) is the mass, \(c\) is the specific heat capacity and \(\Delta T\) is the difference in temperature expressed as \(\Delta T=T_{2}-T_{1}\) You may assume that the specific heat capacity remains constant in your temperature interval, with a value \[\Large c _{water}=4.18 \sf \frac{ J }{ g \times K }\] Solve for mass, \(m\), and evaluate with the given data,